# Prompt-Based Document Modifications In Ranking Competitions

### v1

@@ Abstract.
We study prompting-based approaches with Large Language Models (LLMs) for modifying documents so as to promote their ranking in a competitive search setting. Our methods are inspired by prior work on leveraging LLMs as rankers. We evaluate our approach by deploying it as a bot in previous ranking competitions and in competitions we organized. Our findings demonstrate that our approach effectively improves document ranking while preserving high levels of faithfulness to the original content and maintaining overall document quality.
@@ 1. Introduction
We present novel ranking incentivized document modification methods for ad hoc retrieval that are based on prompting large language models (LLMs). The desiderata of the modifications are (cf., [7]): (i) potentially improving future ranking for a query at hand, (ii) maintaining “faithfulness” to the original document, and (iii) producing a document of high (content) quality. Some of our methods are inspired by recent work on using prompting techniques to have LLMs induce rankings in response to queries [12, 19, 13]. For example, some of the prompts include information induced from past rankings to provide the LLM with implicit information about the undisclosed ranking function.
We evaluate the document modification methods using datasets that resulted from past ranking competitions [20, 14]. In addition, we organized ranking competitions and used our modification methods as bots competing against humans [^1] The dataset of the competition we organized, and the accompanying code, will be made public upon publication of this paper. They are available for reviewing purposes at https://github.com/promptdriven2025.. The empirical evaluation demonstrates the merits of the most effective methods we studied with respect to humans and to a highly effective feature-based approach for ranking-incentivized document modification [7].
Related work.  We refer the reader to [11] for a survey on competitive search. The work most related to ours is that of Goren et al. [7] who devised a supervised feature-based ranking-incentivized document modification method. We discuss the method in Section 3 and use it as a baseline.
@@ 5. Conclusions
We presented novel methods for ranking incentivized document modifications: modifying a document so it will be highly ranked by an undisclosed ranking function for a query. Our methods are based on prompting large language models (LLMs). Some of our methods are inspired by prompting approaches for inducing LLM-based ranking over documents.

### v2

# White Hat Search Engine Optimization using Large Language Models

@@ Abstract.
We present novel white-hat search engine optimization techniques based on genAI and demonstrate their empirical merits.
@@ 1. Introduction
We present novel document modification methods for ad hoc retrieval that are based on prompting large language models (LLMs). The desiderata of the modifications are (cf., [7]): (i) potentially improving future ranking for a query at hand, (ii) maintaining “faithfulness” to the original document, and (iii) producing a document of high (content) quality. Some of our methods are inspired by recent work on using prompting techniques to have LLMs induce rankings in response to queries [12, 19, 13]. For example, some of the prompts include information induced from past rankings to provide the LLM with implicit information about the undisclosed ranking function.
We evaluate the document modification methods using datasets that resulted from past ranking competitions [20, 14]. In addition, we organized ranking competitions and used our modification methods as bots competing against humans [^1] The dataset of the competition we organized, and the accompanying code, will be made public upon publication of this paper. They are available for reviewing purposes at https://github.com/promptdriven2025.. The empirical evaluation demonstrates the merits of the most effective methods we studied with respect to humans and to a highly effective feature-based approach for document modification [7].
Related work.  We refer the reader to [11] for a survey on competitive search. The work most related to ours is that of Goren et al. [7] who devised a supervised feature-based document modification method. We discuss the method in Section 3 and use it as a baseline.
@@ 5. Conclusions
We presented novel methods for document modifications: modifying a document so it will be highly ranked by an undisclosed ranking function for a query. Our methods are based on prompting large language models (LLMs). Some of our methods are inspired by prompting approaches for inducing LLM-based ranking over documents.